OpenStudy (anonymous):

3. Kaiden deposited $6200 into a savings account for which interest is compounded weekly at a rate of 2.77%. How much interest will he earn after 8 years? (Points : 1)

OpenStudy (kropot72):

\[Amount\ after\ 8\ years=6200(1+\frac{0.0277}{52})^{(8\times52)}=6200\times(1.000532692)^{416}=you\ can\ calculate\]

OpenStudy (anonymous):

use the formula I=PRT

OpenStudy (kropot72):

The formula is as follows: \[A=P(1+\frac{r}{n})^{nt}\] A is the amount after t years. P is the principal. r is the annual interest rate as a decimal. n is the number of compounding periods per year.